# The disk precession in a Be star-magnetar binary and its application to the rotation measure of FRB 20201124A

### v1

@@ Abstract
Fast radio bursts (FRBs) are bright, millisecond-duration radio bursts with poorly known origins. Most FRB sources are detected only once, while some are repeaters. Variation patterns observed in the rotation measure (RM) of some repeaters — indicate that the local magneto-ionic environments of these FRB sources are highly dynamic. It has been suggested that a Be star-magnetar binary system is a possible origin for such variation. FRB 20201124A is notable among these sources since it is the most active one and exhibits substantial temporal variations of RM measured by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). The physics behind this long-term behavior is poorly understood. Here we propose that, within the framework of the Be star-magnetar binary scenario, the observed variation of RM is attributed to a combination of orbital motion and the precession of the circumstellar disk of the Be star. While a $\sim 785$-day precession of the disk contributes to the observed decrease in the amplitude of the variation, our model predicts that the amplitude oscillates with this period.
@@ 1 Introduction
FRBs are a class of cosmological radio transients with millisecond durations and high luminosity (Lorimer_2007Sci...318.. [^7] L; Thornton_2013Sci...341...53T; Wu2024). Although the origin of FRBs remains poorly understood, the detection of FRB 200428 from the Galactic magnetar SGR J1935+2154 reveals that magnetars are a viable source of FRBs (Bochenek_2020Natur.587...59B; CHIME_2020Natur.587...54C). Many researchers indicate that the FRB emission region is likely to be within the magnetosphere of the magnetar (Kumar_2017MNRAS.468.2726K; Yang2018; Nimmo_2025Natur.637...48N). Approximately 900 FRB sources have been detected to date, most of them are detected only once, whereas others exhibit repeated emissions (Spitler_2016Natur.531..202S; CHIME_2019ApJ...885L..24C; CHIME_2023ApJ...947...83C). Long-term observations of FRB repeaters allow for the investigation of their properties, e.g., dispersion measure (DM), rotation measure (RM), polarization, burst rate, and waiting time, helping to reveal the origin of FRBs.
@@
Over a decade of observations of FRBs, several FRBs exhibit dramatic RM variations and oscillations (Hilmarsson_2021MNRAS.508.5354H; Xu_2022Natur.609..685X; Kumar_2023MNRAS.526.3652K; Anna-Thomas_2023Sci...380..599A; Mckinven_2023ApJ...950...12M; Ng_2025ApJ...982..154N). Since the RM is related to the magnetic field and plasma density, the temporal variation may indicate the evolution of the local magneto-ionic environment around the FRB source.
@@
Among these FRBs, FRB 20201124A attracts significant attention. It is an extremely active repeater and exhibits peculiar variations in its RM. From the observations by FAST between April and June 2021 (Xu_2022Natur.609..685X), the RM of FRB 20201124A exhibits a pattern of rapid increase, sharp drop, and subsequent re-ascent over $\sim$ 40 days, with the maximum RM variation reaching $\sim$ 500 $\rm{rad/m^{2}}$. The variation stage is overlaid by a series of oscillations, and followed by a stable stage lasting for $\sim$ 20 days. After subtracting the RM averaged over $\sim$ two-month observations (from 1 April to 11 June, 2021), the variations show that the local RM contribution of FRB 20201124A changes its sign (Wang_2022NatCo..13.4382W).
@@
After the 16-day periodically modulated activity for FRB 20180916B was found, the orbital motion of binary systems is proposed (Ioka_&_Zhang_2020ApJ...893L..26I; Katz_2020MNRAS.494L..64K; Sridhar_2021ApJ...917...13S; Pleunis_2021ApJ...911L...3P). However, these models cannot account for the frequency-dependent active window (2021Natur.596..505P). For the observed RM variation of FRB 20201124A, Wang_2022NatCo..13.4382W examined a Be star and magnetar binary scenario and suggested that the orbital motion causes a variation in the local magneto-ionic environment along the line of sight (LoS), resulting in the sign-changing RM. This model is supported by the following studies. Xu_2025arXiv250506006X claimed the detection of a $\sim$ 26-day periodicity with high significance in the RM observations of FRB 20201124A from 2021 to 2022. Du_2025arXiv250312013D detected a $\sim$ 1.7 s periodicity with high significance in the times of arrival (TOAs) of FRB 20201124A bursts observed on MJD 59310 and MJD 59347; such a period is explained by the spinning of the potential magnetar central engine of FRB 20201124A. Zhang_2025arXiv250517880Z further notes that the two days with detected TOAs periodicity bookend the RM variation phase of FRB 20201124A, and they might be the two observational windows when the magnetar central engine crossed the Be star disk. In addition to FRB 20201124A, Liang_2025arXiv250510463L reported the existence of a $\sim$ 200-day periodicity in the RM of FRB 20220529 from 2022 to 2025. The RM variation of FRB 20180916B may be caused by the stellar wind of a massive companion (Zhao2023). These potential periodicities in RM are likely to be related to the orbital motion of binary star systems (Rajwade2023; ZhangB2025; Wang2025).
@@
Super-orbital brightness variations — periodic brightness oscillations far longer than the binary’s orbital period in multi-wavelength bands — have been observed in many Be star-neutron star binaries (Ogilvie_2001MNRAS.320..485O; McGowan_2008MNRAS.384..821M; Rajoelimanana_2011MNRAS.413.1600R) — specifically a Be/X-ray binary (Negueruela_1998A&A...336..251N; Coe_2005MNRAS.356..502C; Liu_2005A&A...442.1135L; Haberl_&_Sturm_2016A&A...586A..81H). While the driving mechanism of this periodic behavior is unclear, many studies indicate that it could be induced by Be disk precession (Rajoelimanana_2011MNRAS.413.1600R; Martin_&_Franchini_2021ApJ...922L..37M; Martin_2023MNRAS.523L..75M; Martin_&_Charles_2024MNRAS.528L..59M; Chen_2024ApJ...973..162C). Assuming that the rotation axis of the Be star is precessing with a superorbital period, and according to Martin_2023MNRAS.523L..75M, the decretion disk can remain locked to the equatorial plane of the Be star; the node of the disk plane and the neutron star orbit will precess with the star. Since the brightness variations of the Be star essentially reflect changes in the disk density along the LoS between the Be star and the observer, the disk precession can lead to the periodic modulation of the density profile along the line of sight, producing the super-orbital modulation in the observed brightness (Alcock_2001MNRAS.321..678A; Rajoelimanana_2011MNRAS.413.1600R; Martin_&_Charles_2024MNRAS.528L..59M). Moreover, X-ray bursts emitted when the neutron star interacts with the disk may also have their brightness modulated by the precession of the Be star and the disk (Alcock_2001MNRAS.321..678A). The magnitude of FRB 20201124A RM varies significantly, which cannot be explained by a simple binary orbital motion. The above studies motivate us to investigate the precession of the Be star disk in binary systems modulating the RM variations of FRBs.
@@
In this work, we present a Be star-magnetar model with a precessing disk to elucidate the long-term RM variations in FRB 20201124A. The paper is organized as follows. The observations and data reduction of FRB 20201124A are presented in Section 2. The model of a binary system with a precessing disk is described in Section 3. The fitting results are shown in Section 4. In Section 5, we draw conclusions and discuss other FRB sources with similar RM variations.
2 FRB 20201124A
@@
FRB 20201124A is a highly active repeating FRB source. By April 2021, FRB 20201124A entered an extremely active phase (Xu_2022Natur.609..685X), making it one of the most active known FRBs. In addition, FRB 20201124A is found to be associated with a persistent radio source (PRS), which supports the magnetar origin of this FRB (Bruni_2024Natur.632.1014B). Since 2021, FAST has been monitoring this source, covering the 1.0–1.5 GHz frequency range, and has detected thousands of bursts during the three active cycles (Xu_2022Natur.609..685X; Jiang_2022RAA....22l4003J; Niu_2022RAA....22l4004N; Xu_2025arXiv250506006X). The three FAST observation epochs are hereinafter referred to as E1, E2, and E3 for brevity. Only the bursts with high signal-to-noise ratio $S/N$ are adopted to measure the RMs by using a Q-U fitting method (Desvignes_2019Sci...365.1013D).
@@
In E1, FAST monitored FRB 20201124A from April 1, 2021, to June 11, 2021 (UT). During a total observing time of 91 hours, 1,863 bursts were detected. A total of 1,103 bright bursts with a $S/N>30$ are adopted to measure the RM (Xu_2022Natur.609..685X). Intriguingly, the RM of the adopted bursts shows dramatic variation in a short term ($\sim 60$ days), and the maximum variation can reach about 500 $\rm{rad\,m^{-2}}$, making it the first FRB observed with rapid RM variation in such a short timescale (Xu_2022Natur.609..685X; Wang_2022NatCo..13.4382W).
@@
E2 spans from 25 September 2021 to 17 October 2021 (UT), with a total observation of 19 hours. However, a total of 881 bursts was observed during the first four days, i.e., 25-28 September 2021, whereas no bursts were detected later in this epoch (Zhou_2022RAA....22l4001Z; Zhang_2022RAA....22l4002Z; Jiang_2022RAA....22l4003J; Niu_2022RAA....22l4004N). Among these 881 bursts, a total of 536 bursts with $S/N\geq 50$ are adopted to measure RM. The RM of the E2 bursts remains at the average level of the E1 bursts, but displays a feature of continuous drop. swinging and orthogonal jumping PAs are also found in some of these bursts (Niu_2024ApJ...972L..20N).
@@
E3 is conducted by FAST from 2 February 2022 to 26 May 2022, during the third active cycle. In this epoch, 1,467 bright bursts with $S/N>20$ are detected in the total observing time of 34.3 hours (Xu_2025arXiv250506006X). The RM data of the E3 bursts also show a variation feature within several tens of days, but at a level of $\sim 100\,\rm{rad\,m^{-2}}$, which is a few times lower than that of epoch one. Xu_2025arXiv250506006X searched for the periodicity of the RM variation in the E3 bursts, and a period of $\sim 30$ days is found with high significance. We note that the E3 RM data have a lower background level compared to that of E1 and E2, which might be due to the long-term evolution of the magneto-ionic environment.
@@
In this work, we use the FRB 20201124A RM data of E1, E2, and E3 (denoted as $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$ below, respectively). The data of $\rm{RM_{E1}}$ are provided by (Xu_2022Natur.609..685X), and the data of $\rm{RM_{E2}}$ and $\rm{RM_{E3}}$ are obtained from the plots in Xu_2025arXiv250506006X using the online data extraction tool WebPlotDigitizer (Marin_2017sf2a.conf..113M). To characterize RM variations, we perform some pre-processing operations on $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$. For each episode (episode $i$), we first bin $\rm{RM_{E\textit{i}}}$ by MJD and derive the mean RM value in each MJD bin (after this step, the data of episode $i$ are then denoted as $\rm{RM_{dE\textit{i}}}$). Secondly, we derive the mean for $\rm{RM_{dE\textit{i}}}$, which gives $\overline{\rm{RM}}_{\rm{dE1}}=-583.95\,\rm{rad}\,m^{-2}$, $\overline{\rm{RM}}_{\rm{dE2}}=-593.35\,\rm{rad}\,m^{-2}$, and $\overline{\rm{RM}}_{\rm{dE3}}=-666.45\,\rm{rad}\,m^{-2}$. Finally, we can obtain the RM variations $\Delta\rm{RM}_{E1}$, $\Delta\rm{RM}_{E2}$, and $\Delta\rm{RM}_{E3}$ by calculating $\Delta\rm{RM}_{E\textit{i}}=\rm{RM_{dE\textit{i}}}-\overline{\rm{RM}}_{dE\textit{i}}$.
3 Be star-magnetar binary with disk precession
3.1 Be star-magnetar binary
@@
We consider a Be star-magnetar binary system. The stellar outflow from the Be star can be particularly significant in the equatorial region, where the dense flow usually forms a decretion disk. The density of the decretion disk decreases with increasing radius, and the base density $\rho_{0}$ typically ranges from about $10^{-12}$ to $10^{-10}\,\rm{g\,cm^{-3}}$ (Porter_1999A&A...348..512P; Gies_2007ApJ...654..527G; Rivinius_2013A&ARv..21...69R; Wang_2022NatCo..13.4382W).
@@
Here, we consider an axisymmetric disk and describe the density profile as (Wang_2022NatCo..13.4382W)
@@
where $r$, $D$, $R_{\star}$, and $\beta$ are the distances from the stellar center, the heights from the stellar equatorial plane, the stellar radius, and the density slope index, respectively. In the case of a Be star, $\beta$ usually ranges from $2\sim 4$. In addition, $H(r)$ is the vertical scale height of the disk, which is determined by (Bjorkman_1997LNP...497..239B; Rivinius_2013A&ARv..21...69R)
@@
Previous studies suggest that a magnetic field exists in the circumstellar disk (Reiners_&Mohanty_2012ApJ...746...43R; Pudritz_2014IAUS..302...10P; Tsukamoto_2016PASA...33...10T; Ohashi_2025NatAs...9..526O). Here, we adopt a simplified toroidal magnetic field (Johnston_1996MNRAS.279.1026J; Brauer_2017A&A...607A.104B)
@@
where $\rm{B_{0}}$, r, $\vec{e}_{r}$, and $\vec{e}_{\rm{disk}}$ are the magnetic field strengths at the stellar surface, the vector pointing from the stellar center to the field point, the unit vector of r, and the unit normal vector of the Be star disk, respectively.
@@
where $g_{\rm H^{+}}$ and $g_{\rm H}$ are the statistical weights of the proton and neutral hydrogen atom, respectively (we set $g_{\rm H^{+}}=g_{\rm H}$ for simplification), $n_{\rm H^{+}}$ is the number density of neutral hydrogen atoms, and we assume that $n_{\rm H^{+}}=n_{\rm e}$ according to the conservation of charge. In addition, $m_{\rm e}$, $h$, and $\chi_{0}=13.6\,\rm{eV}$ are the electron mass, Planck constant, and ionization energy of the hydrogen atom, respectively.
Figure 1: A schematic diagram of the binary system in Section 3 (not to scale). Note that the darker regions in the disk have higher density. The magnetar continuously emits FRBs and orbits the Be star along an elliptical orbit. FRBs emitted by the magnetar at different positions along its orbit propagate along the LoS and reach the observer. The precession of the Be star disk leads to varying RMs among different FRBs.
3.2 Disk precession
In a Be star-NS binary, the supernova natal kick at the birth of the NS may lead to the misalignment of the orbital plane and the Be star disk plane (Brandt_1995MNRAS.274..461B),as well as a binary orbit with high eccentricity (van_den_Heuvel_1997ApJ...483..399V).
Following the method of Chen_2024ApJ...973..162C, we set a Cartesian coordinate system where the origin is at the center of the star, the $x$-axis points to the periastron of the binary star orbit, and the $z$-axis is perpendicular to the orbital plane (see Figure 1). The unit vectors of the $x$-axis, $y$-axis and $z$-axis, are denoted by $\vec{e}_{x}$, $\vec{e}_{y}$ and $\vec{e}_{z}$, respectively. The unit direction vector of the magnetar $\vec{e}_{\rm{mag}}$, and the angle between it and $x$ axis is $\psi$. The relation between the true anomaly $\psi$ and the time $t$ (orbital phase) is described by Kepler’s equation:

### v2

@@ Abstract
Fast radio bursts (FRBs) are bright, millisecond-duration radio bursts with poorly known origins. Most FRB sources are detected only once, while some are repeaters. Variation patterns observed in the rotation measure (RM) of some repeaters indicate that the local magneto-ionic environments of these FRB sources are highly dynamic. It has been suggested that a Be star-magnetar binary system is a possible origin for such variation. FRB 20201124A is notable among these sources since it is one of the very active repeaters and exhibits substantial temporal variations of RM measured by the Five-hundred-meter Aperture Spherical radio Telescope (FAST). The physics behind this long-term behavior remains poorly understood. Here we propose that, within the framework of the Be star-magnetar binary scenario, the observed variation in RM is attributed to a combination of orbital motion and the precession of the circumstellar disk of the Be star. While a $\sim 785$-day precession of the disk contributes to the observed decrease in the amplitude of the variation, our model predicts that the amplitude oscillates with this period.
@@ I Introduction
FRBs are a class of cosmological radio transients with millisecond durations and high luminosity (Lorimer et al., 2007; Thornton et al., 2013; Wu and Wang, 2024). Although the origin of FRBs remains poorly understood, the detection of FRB 200428 from the Galactic magnetar SGR J1935+2154 reveals that magnetars are a viable source of FRBs (Bochenek et al., 2020; CHIME/FRB Collaboration et al., 2020). Many researchers indicate that FRB emission regions are likely to be within magnetar magnetospheres (Kumar et al., 2017; Yang and Zhang, 2018; Nimmo et al., 2025). Approximately 4,000 FRB sources have been detected to date (FRB Collaboration et al., 2026), most of which are detected only once, whereas others exhibit repeated emissions (Spitler et al., 2016; CHIME/FRB Collaboration et al., 2019; Chime/Frb Collaboration et al., 2023). Long-term observations of FRB repeaters allow for the investigation of their properties, e.g., dispersion measure (DM), rotation measure (RM), polarization, burst rate, and waiting time, helping to reveal the origin of FRBs.
@@
Over a decade of observations of FRBs, several FRBs have exhibited dramatic RM variations and oscillations (Hilmarsson et al., 2021; Xu et al., 2022; Kumar et al., 2023; Anna-Thomas et al., 2023; Mckinven et al., 2023; Ng et al., 2025). Since the RM is related to the magnetic field and plasma density, the temporal variation may indicate the evolution of the local magneto-ionic environment around the FRB source.
@@
Among these FRBs, FRB 20201124A attracts significant attention. It is an extremely active repeater and exhibits peculiar variations in its RM. From the observations by FAST between April and June 2021 (Xu et al., 2022), the RM of FRB 20201124A exhibits a pattern of rapid increase, sharp drop, and subsequent re-ascent over $\sim$ 40 days, with the maximum RM variation reaching $\sim$ 500 $\rm{rad/m^{2}}$. The variation stage is overlaid by a series of oscillations, and followed by a stable stage lasting for $\sim$ 20 days. After subtracting the RM averaged over $\sim$ two-month observations (from 1 April to 11 June, 2021), the variations show that the local RM contribution of FRB 20201124A changes its sign (Wang et al., 2022).
To explain the origin of repeating FRBs, a model involving the interaction between a pulsar magnetosphere and its companion star’s stream, i.e., the ”cosmic-comb” model, has been proposed by Zhang (2017). Later this model is adopted in Zhang (2018) to explain the observational properties of FRB 20121102. After the 16-day periodically modulated activity for FRB 20180916B is found, more studies focused on the relation between FRBs and the orbital motion of binary systems are conducted (Ioka and Zhang, 2020; Katz, 2020; Sridhar et al., 2021; Pleunis et al., 2021)  [^1] These models cannot account for the observed frequency-dependent active window of FRB 20180916B (Pastor-Marazuela et al., 2021), until Wada et al. (2021) offers an explanation via the “cosmic-comb” model..
@@
For the observed RM variation of FRB 20201124A, Wang et al. (2022) examined a Be star and magnetar binary scenario and suggested that the orbital motion causes a variation in the local magneto-ionic environment along the line of sight (LoS), resulting in the sign-changing RM. This model is supported by the following studies (Xu et al., 2025). Zhao et al. (2023) points out that the RM variation of FRB 20180916B may be caused by the stellar wind of a massive companion. Xu et al. (2025) claimed the detection of a $\sim$ 26-day periodicity with significance $5\sigma$ in the RM observations of FRB 20201124A from 2021 to 2022. Du et al. (2025) detected a $\sim$ 1.7 s periodicity with high significance in the times of arrival (ToAs) of FRB 20201124A bursts observed on MJD 59310 and MJD 59347; such a period is explained by the spinning of the potential magnetar central engine of FRB 20201124A. Zhang et al. (2025) further noted that the two days with detected ToAs periodicity bookend the RM variation phase of FRB 20201124A, and they might be the two observational windows when the magnetar central engine crossed the Be star disk. In addition to FRB 20201124A, Li et al. (2026) reported that the RM of FRB 20220529 shows sudden variations within a month, and Liang et al. (2025) reported the existence of a $\sim$ 200-day periodicity in the RM of FRB 20220529 from 2022 to 2025. These potential periodicities in RM are likely to be related to the orbital motion of binary star systems (Rajwade and van den Eijnden, 2023; Zhang and Hu, 2025; Wang et al., 2025).
@@
Super-orbital brightness variations — periodic brightness oscillations far longer than the binary’s orbital period in multi-wavelength bands — have been observed in many Be star-neutron star binaries (Ogilvie and Dubus, 2001; McGowan et al., 2008; Rajoelimanana et al., 2011) — specifically a Be/X-ray binary (Negueruela et al., 1998; Coe et al., 2005; Liu et al., 2005; Haberl and Sturm, 2016). While the driving mechanism of this periodic behavior is unclear, many studies indicate that it could be induced by Be disk precession (Rajoelimanana et al., 2011; Martin and Franchini, 2021; Martin, 2023; Martin and Charles, 2024; Chen et al., 2024). Assuming that the rotation axis of the Be star precesses with a superorbital period, and according to Martin (2023), the decretion disk can remain locked to the equatorial plane of the Be star; the node of the disk plane and the neutron star orbit will precess with the star. Since the brightness variations of the Be star essentially reflect changes in the disk density along the LoS between the Be star and the observer, the disk precession can lead to the periodic modulation of the density profile along the line of sight, producing the super-orbital modulation in the observed brightness (Alcock et al., 2001; Rajoelimanana et al., 2011; Martin and Charles, 2024). Moreover, X-ray bursts emitted when the neutron star interacts with the disk may also have their brightness modulated by the precession of the Be star and the disk (Alcock et al., 2001). The magnitude of FRB 20201124A RM varies significantly in three active epochs, which cannot be explained by simple binary orbital motion. The above studies motivate us to investigate the precession of the Be star disk in binary systems and its effect on modulating the RM variations of FRBs.
@@
In this work, we present a Be star-magnetar model with a precessing disk to elucidate the long-term RM variations in FRB 20201124A. The paper is organized as follows. The observations and data reduction of FRB 20201124A are presented in Section II. The model of a binary system with a precessing disk is described in Section III. The fitting results are shown in Section IV. In Section V, we draw conclusions and discuss other FRB sources with similar RM variations.
II FRB 20201124A
@@
FRB 20201124A is a highly active repeating FRB source. By April 2021, FRB 20201124A entered an extremely active phase (Xu et al., 2022). In addition, FRB 20201124A is found to be associated with a persistent radio source (PRS), which is consistent with the magnetar origin hypothesis for this FRB (Bruni et al., 2024). Since 2021, FAST has been monitoring this source, covering the 1.0–1.5 GHz frequency range, and has detected thousands of bursts during the three active cycles (Xu et al., 2022; Jiang et al., 2022; Niu et al., 2022; Xu et al., 2025). The three FAST observation epochs are hereinafter referred to as E1, E2, and E3 for brevity. Only the bursts with a high signal-to-noise ratio $S/N$ are adopted to measure the RMs using a Q-U fitting method (Desvignes et al., 2019).
@@
In E1, FAST monitored FRB 20201124A from April 1, 2021, to June 11, 2021 (UT). During a total observing time of 91 hours, 1,863 bursts were detected. A total of 1,103 bright bursts with a $S/N>30$ are adopted to measure the RM (Xu et al., 2022). Intriguingly, the RM of the adopted bursts shows dramatic variation in a short term ($\sim 60$ days), and the maximum variation can reach about 500 $\rm{rad\,m^{-2}}$, making it the first FRB observed with rapid RM variation in such a short timescale (Xu et al., 2022; Wang et al., 2022).
@@
E2 spans from 25 September 2021 to 17 October 2021 (UT), with a total observation of 19 hours. However, a total of 881 bursts was observed during the first four days, i.e., 25-28 September 2021, whereas no bursts were detected later in this epoch (Zhou et al., 2022; Zhang et al., 2022; Jiang et al., 2022; Niu et al., 2022). Among these 881 bursts, 536 bursts with $S/N\geq 50$ are adopted to measure RM. The RM of the E2 bursts remains at the average level of the E1 bursts, but displays a feature of continuous drop. swinging and orthogonal jumping PAs are also found in some of these bursts (Niu et al., 2024).
@@
E3 was conducted by FAST from 2 February 2022 to 26 May 2022, during the third active cycle. In this epoch, 1,467 bright bursts with $S/N>20$ are detected in a total observing time of 34.3 hours (Xu et al., 2025). The RM data of the E3 bursts also show a variation feature within several tens of days, but at a level of $\sim 100\,\rm{rad\,m^{-2}}$, which is a few times lower than that of epoch one. Xu et al. (2025) searched for the periodicity of the RM variation in the E3 bursts, and a period of $\sim 30$ days was found with high significance. We note that the E3 RM data have a lower background level compared to that of E1 and E2, which might be due to the long-term evolution of the magneto-ionic environment. After E3, this source falls into quiescence.
@@
In this work, we use the FRB 20201124A RM data of E1, E2, and E3 (denoted as $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$ below, respectively). Data of $\rm{RM_{E1}}$ are provided by (Xu et al., 2022), and data of $\rm{RM_{E2}}$ and $\rm{RM_{E3}}$ are obtained from the plots in Xu et al. (2025) using the online data extraction tool WebPlotDigitizer (Marin et al., 2017). To characterize RM variations, we perform some pre-processing operations on $\rm{RM_{E1}}$, $\rm{RM_{E2}}$, and $\rm{RM_{E3}}$. For each episode (episode $i$), we first bin $\rm{RM_{E\textit{i}}}$ by MJD and derive the mean RM value in each MJD bin (after this step, the data of episode $i$ are then denoted as $\rm{RM_{dE\textit{i}}}$). Secondly, we derive the mean for $\rm{RM_{dE\textit{i}}}$, which gives $\overline{\rm{RM}}_{\rm{dE1}}=-583.95\,\rm{rad}\,m^{-2}$, $\overline{\rm{RM}}_{\rm{dE2}}=-593.35\,\rm{rad}\,m^{-2}$, and $\overline{\rm{RM}}_{\rm{dE3}}=-666.45\,\rm{rad}\,m^{-2}$. Finally, we can obtain the RM variations $\Delta\rm{RM}_{E1}$, $\Delta\rm{RM}_{E2}$, and $\Delta\rm{RM}_{E3}$ by calculating $\Delta\rm{RM}_{E\textit{i}}=\rm{RM_{dE\textit{i}}}-\overline{\rm{RM}}_{dE\textit{i}}$.
III Be star-magnetar binary with disk precession
III.1 Be star-magnetar binary
@@
We consider a Be star-magnetar binary system. Stellar outflow from the Be star can be particularly significant in the equatorial region, where dense flow usually forms a decretion disk. The density of the decretion disk decreases with increasing radius, and the base density $\rho_{0}$ typically ranges from about $10^{-12}$ to $10^{-10}\,\rm{g\,cm^{-3}}$ (Porter, 1999; Gies et al., 2007; Rivinius et al., 2013; Wang et al., 2022).
@@
Here, we consider an axisymmetric disk and describe the density profile as (Wang et al., 2022)
@@
where $\rho_{0}$, $r$, $D$, $R_{\star}$, and $\beta$ are the disk density at the stellar radius, the distances from the stellar center, the heights from the stellar equatorial plane, the stellar radius, and the density slope index, respectively. In the case of a Be star, $\beta$ usually ranges from $2\sim 4$. In addition, $H(r)$ is the vertical scale height of the disk, which is determined by (Bjorkman, 1997; Rivinius et al., 2013)
@@
Previous studies indicate a magnetic field is present in the outflow disk of a early type star like Be star (Usov and Melrose, 1992; Melatos et al., 1995; Wang et al., 2022; Zhao et al., 2023). Here, we adopt a simplified toroidal magnetic field at a relatively large distance from the Be star:
@@
where $\rm{B_{0}}$, r, $\vec{e}_{r}$, and $\vec{e}_{\rm{disk}}$ are the initial strength of the toroidal magnetic field, the vector pointing from the stellar center to the field point, the unit vector of r, and the unit normal vector of the Be star disk, respectively. A more detailed discussion on the magnetic field configuration of Be stars can be seen in Section V.1.4.
@@
where $g_{\rm H^{+}}$ and $g_{\rm H}$ are the statistical weights of the proton and neutral hydrogen atom, respectively (we set $g_{\rm H^{+}}=g_{\rm H}$ for simplification), $n_{\rm H^{+}}$ is the number density of ionized hydrogen atoms, and we assume that $n_{\rm H^{+}}=n_{\rm e}$ according to the conservation of charge. In addition, $m_{\rm e}$, $h$, and $\chi_{0}=13.6\,\rm{eV}$ are the electron mass, Planck constant, and ionization energy of the hydrogen atom, respectively.
Figure 1: A schematic diagram of the binary system in Section III (not to scale). Note that the darker regions in the disk have higher density. The magnetar continuously emits FRBs and orbits the Be star along an elliptical orbit. FRBs emitted by the magnetar at different positions along its orbit propagate along the LoS and reach the observer. Precession of the Be star disk leads to varying RMs among different bursts.
III.2 Disk precession
In a Be star-NS binary, the supernova natal kick at the birth of the NS may lead to a misalignment of the orbital plane and the Be star disk plane (Brandt and Podsiadlowski, 1995),as well as a binary orbit with high eccentricity (van den Heuvel and van Paradijs, 1997).
Following the method of Chen et al. (2024), we set a Cartesian coordinate system where the origin is at the center of the star, where the $x$-axis points toward the periastron of the binary star orbit, and the $z$-axis is perpendicular to the orbital plane (see Figure 1). The unit vectors of the $x$-, $y$- and $z$-axes, are denoted by $\vec{e}_{x}$, $\vec{e}_{y}$ and $\vec{e}_{z}$, respectively. The unit direction vector of the magnetar $\vec{e}_{\rm{mag}}$, and the angle between it and $x$ axis is $\psi$. The relation between the true anomaly $\psi$ and the time $t$ (orbital phase) is described by Kepler’s equation: